Water is the working fluid in a Rankine cycle. Superheated vapor enters the turbine at 10 Mpa, 560 C with a mass flow rate of 7.8kg/s and exits at 8 kPa. Saturated liquid enters the pump at 8 kPa. The isentropic turbine efficiency is 85%, and the isentropic pump efficiency is 85%. Cooling water enters the adiabatic condenser at 18 C and exits at 36 C with no significant change in pressure and assuming the specific heat of the cooling water remains constant at cp = 4.179 kJ/kg-K.

a.) Enter the enthalpy values at all state points.

b.) Find the net power produced (kW)

c.) Find the heat transferred from the boiler to the steam (kW)

d.) Find the thermal efficiency of the cycle.

e.) Heat transferred from the steam in the condenser (kW)

f.) Mass flow rate of condenser cooling water (kg/s)
Answer #1

- Ruperheated steam enters h furbine t 10M Pa、530 ,c. seturted (it enten-tte. Pump at SE:Pa 予 Mno).turbin.sgs72%) pum h,-35203520.-180598 3333,462-Ble,K7, 8 =26047, 33 kis (Yout -2-h3 =よ318-53-16875-2143.78 = (6768-2tyEQ (η =m(p)&T 16768.28 4 rio: 16